# ASCoT: An Adaptive Self-Correction Chain-of-Thought Method for Late-Stage Fragility in LLMs

### v3

@@ Abstract.
Chain-of-Thought (CoT) prompting has significantly advanced the reasoning capabilities of Large Language Models (LLMs), yet the reliability of these reasoning chains remains a critical challenge. A widely held “cascading failure” hypothesis suggests that errors are most detrimental when they occur early in the reasoning process. This paper challenges that assumption through systematic error-injection experiments, revealing a counter-intuitive phenomenon we term “Late-Stage Fragility”: errors introduced in the later stages of a CoT chain are significantly more likely to corrupt the final answer than identical errors made at the beginning. To address this specific vulnerability, we introduce the Adaptive Self-Correction Chain-of-Thought (ASCoT) method. ASCoT employs a modular pipeline in which an Adaptive Verification Manager (AVM) operates first, followed by the Multi-Perspective Self-Correction Engine (MSCE). The AVM leverages a Positional Impact Score function that assigns different weights based on the position within the reasoning chains, addressing the Late-Stage Fragility issue by identifying and prioritizing high-risk, late-stage steps. Once these critical steps are identified, the MSCE applies robust, dual-path correction specifically to the failure parts. Extensive experiments on benchmarks demonstrate our superiority over previous methods. ASCoT reduces token usage by 21%/30% for the LLaMA-3.1-8B on GSM8K/MATH-500 benchmarks, with minimal accuracy drops of only 1.1% and 1.8%, achieving new sota performance.
Key words and phrases:
Large Language Models, Common-Sense Reasoning, Interpretability and Evaluation of NLP Models, Prompt Engineering
1. Introduction
@@
Figure 1. The top part (a) illustrates the standard CoT reasoning process, which often generates redundant outputs and lacks an effective self-correction mechanism when errors occur. The bottom part (b) depicts the ASCoT method, which incorporates a robust mechanism for self-correction, addressing potential reasoning flaws at each stage.
@@
The advent of Chain-of-Thought (CoT) has endowed Large Language Models (LLMs) with remarkable multi-step reasoning capabilities Kojima et al. (2022), enabling them to deconstruct complex problems into a sequence of manageable sub-problems Yao et al. (2023); Shinn et al. (2023). Recent studies, such as the o1 from OpenAI Jaech et al. (2024) and DeepSeekR1 Guo et al. (2025) have shown that extended CoT reasoning improves LLM performance. However, the success of CoT and its variants has introduced challenges: prohibitive computational costs from long reasoning chains Chen et al. (2025) and the inherent brittleness of the reasoning process itself Zhang et al. (2024); Ke et al. (2025). A single flaw is often sufficient to invalidate an entire process of thought, raising significant concerns about the trustworthiness of these models Chowdhury and Caragea (2025).
@@
A widely held belief within the research community, often termed the “cascading failure” hypothesis, posits that the primary threat to CoT integrity lies in early-stage errors Tyen et al. (2023). The logic is intuitive: an initial mistake in problem comprehension or a miscalculation will inevitably propagate through the logical dependency graph of the reasoning chain, contaminating all subsequent steps. In this work, we directly investigate this foundational assumption, seeking to answer a critical question: Are all errors in a CoT created equal in their impact? To find out this, we conduct a systematic error-injection experiment, which introduces controlled symbolic and numeric errors at varying positions within correct reasoning chains. We then instruct the LLM to resume the reasoning process from the point at which the error occurred. Our findings are striking and counter-intuitive. We discover that errors introduced in later stages are more likely to lead to an incorrect final answer. We term this phenomenon “Late-Stage Fragility”. An early error often appears to trigger the model’s latent self-correction mechanisms, allowing it to recover and still reach the correct solution.
@@
As the model progresses steadily, its reasoning process gains a form of “semantic commitment”, reducing relevant information entropy. At the later stages, the model is unable to evaluate the validity of the final calculations. This empirical discovery suggests that a significant portion of efforts to secure CoT reasoning may be misdirected. As illustrated in Figure 1(a), a standard CoT process can be both inefficient due to redundancy and unreliable due to its inability to correct these late-stage errors. Building on a rethinking of how we evaluate and how we correct LLM reasoning, we introduce the ASCoT method. In contrast to standard methods, ASCoT is designed to address both challenges simultaneously. As shown in Figure 1(b), it first employs a semantic pruning stage to enhance efficiency, followed by a targeted, risk-aware correction mechanism that focuses computational resources precisely on the most vulnerable steps to ensure a correct final answer.
To sum up, our contributions are summarized as follows:
@@
  • As far as we know, this is the first to identify and quantify the “Late-Stage Fragility” phenomenon in CoT reasoning, demonstrating that errors in later steps are significantly more detrimental than those in earlier steps.
@@
  • We introduce ASCoT, a novel method specifically designed to counteract this fragility by precisely identifying and robustly correcting the high-impact, late-stage errors that are most critical to reasoning accuracy.
@@
  • Our extensive experiments validate the effectiveness of ASCoT. We demonstrate that ASCoT achieves outstanding accuracy on the GSM8k and MATH benchmarks, validating our approach of diagnosing the specific failure mode before engineering a targeted solution.
2. Preliminary
@@
This section establishes the foundational concepts that motivate our work, offering a review of some effective methods for identifying token importance within CoT.
@@
Efficient CoT in LLM Reasoning A critical question for CoT efficiency is whether every token in different position contributes equally to the final answer Xia et al. (2025). Several early approaches Hou et al. (2022); Zhang et al. (2023); Chen et al. (2024b) measure importance based on the self-information or surprisal from a unidirectional language model:
@@
To address this, more advanced methods like LLMLingua-2 Pan et al. (2024) utilize a bidirectional language model Jawahar et al. (2019); Devlin et al. (2019) (e.g., BERT-style) to reframe the task as token classification. In this paradigm, the importance of a token $x_{i}$ is defined as the probability that a trained classifier assigns it an ”essential” label ($y_{i}=1$), conditioned on the entire sequence $x$:
@@
where $\theta_{\mathcal{M}{cls}}$ represents the parameters of the trained bidirectional classifier. Such methods confirm that tokens within a CoT sequence are variable in importance. For example, tokens comprising mathematical equations or key entities Ma et al. (2025) are identified as critical, whereas semantic connectors such as so” and since” contribute less to the final outcome of the reasoning Hou et al. (2025). This highlights that CoT efficiency can be improved by pruning these redundant tokens.
Error Propagation in Reasoning Process Beyond efficiency, the primary challenge for CoT is reliability. A single error can invalidate an entire reasoning process Guo et al. (2017); Tyen et al. (2023). The conventional understanding of this problem is rooted in the “cascading failure” hypothesis, which posits that errors are most damaging when they occur early in the reasoning chain Jiang et al. (2025). According to this view, an early mistake propagates through the logical dependency graph, contaminating all subsequent steps. In this study, through systematic error-injection experiments, we observe that errors introduced in the later stages of a reasoning chain are significantly more detrimental to the final outcome than identical errors introduced at the beginning.
Figure 2. Illustration of ASCoT Pipeline. ASCoT generates the $CoT_{initial}$ from the target LLM, which is then compressed by IRM module to a predefined ratio $\gamma$. The compressed CoT is verified by AVM module, and if the confidence score exceeds a threshold $\tau$, the problematic steps are sent to MSCE module for error correction. The corrected $CoT_{final}$ is then used for fine-tuning. Finally, ASCoT enables reasoning on new problems with the specified compression ratio $\gamma$.
Algorithm 1 ASCoT Mechanism
Input: Initial chain-of-thought (CoT) $CoT_{initial}$
Output: Corrected $CoT_{final}$
[…]
3:for each step $t_{k}$ in $CoT_{pruned}$ do
4:  $is\_flagged\leftarrow\mathrm{AVM}(context,t_{k})$
5:  if is_flaggedis\_flagged then
6:   $t_{k}^{\mathrm{corr}}\leftarrow\mathrm{MSCE}(context,t_{k})$
7:   append $t_{k}^{\mathrm{corr}}$ to $CoT_{final}$
8:   $context\leftarrow context\,\|\,t_{k}^{\mathrm{corr}}$
9:  else
10:   append $t_{k}$ to $CoT_{final}$
11:   $context\leftarrow context\,\|\,t_{k}$
12:  end if
13:end for
14:return $CoT_{final}$
3. Method
In this section, we detail the architecture of our method. ASCoT generates the $CoT_{initial}$ based on the target LLM. The $CoT_{initial}$ is then processed through the IRM module, reducing the output to a fixed ratio $\gamma$. The remaining CoT is passed through the AVM module, which performs a rigorous verification process. If the model’s risk score for any CoT step exceeds a predefined threshold $\tau$, indicating a potential error, the corresponding CoT step is forwarded to the MSCE module. The MSCE module applies internal and external validation techniques to correct errors, resulting in the corrected $CoT_{final}$. Ultimately, the refined, non-redundant $CoT_{final}$ is used for fine-tuning. We illustrate the training and inference process in Figure 2, respectively. We also provide Pseudocode 1 for ASCoT.
3.1. Intelligent Routing Mechanism
@@
3.2. Adaptive Verification Manager
@@
Figure 3. The details of verification and self-correction. AVM and MSCE working together to correct a faulty step $t_{k}$. AVM computes $R(t_{k})$. MSCE then applies intrinsic and extrinsic correction to produce the final corrected $CoT_{final}$.
Positional Impact Score To quantitatively model the “Late-Stage Fragility” phenomenon observed in our error-injection experiments (Section 4.3), we introduce the Position Impact Function $I(k)$. This function is an empirical model fitted to our experimental data, designed to capture the escalating impact of errors at later reasoning stages. We normalize this position function as:
@@
where $k$ is the step’s index, $K$ is the total number of steps, $w_{a}$ is a base impact coefficient, and $\alpha$ is a propagation rate. These parameters are determined by fitting the function to our error-injection data. This procedure involves injecting single errors at different steps $k$ in a set of baseline correct reasoning chains and measuring the probability of final answer failure. The function $I(k)$ is then fitted to these empirical data points. Details of the analysis are available in Appendix.
@@
This score is then compared against a tunable system threshold $\tau$. If $R(t_{k})>\tau$, the step is flagged as high-risk and passed to the MSCE for correction.
Table 1. Experimental results of ASCoT on LLaMA-3.1-8B-Instruct. We report Accuracy, average CoT token count (Tokens) and actual compression ratio (ActRatio) for comparison. The compression ratio $\gamma$ is set to 0.51.00.5-1.0, respectively.
@@
Original 1.0 86.2(0.0↓) 213.17 1.00 48.6(0.0↓) 502.60 1.00
Prompt 0.9 84.1(2.1↓) 226.37 1.06 48.6(0.0↓) 468.04 0.93
0.7 84.9(1.3↓) 209.39 0.98 48.4(0.4↓) 472.13 0.94
0.5 83.7(2.5↓) 188.82 0.89 47.8(0.4↓) 471.11 0.94
Truncation 0.9 70.2(26.0↓) 202.06 0.95 47.8(0.8↓) 440.33 0.88
0.7 25.9(60.3↓) 149.97 0.70 45.0(3.6↓) 386.89 0.77
0.5 7.0(79.2↓) 103.69 0.49 27.4(21.2↓) 283.70 0.56
ASCoT 1.0 86.9(0.7↑) 214.24 1.00 48.8(0.2↑) 478.80 0.95
0.9 86.1(0.1↓) 197.65 0.92 47.6(1.0↓) 427.15 0.84
0.8 85.1(1.1↓) 169.96 0.79 47.2(1.4↓) 376.72 0.74
0.7 83.7(2.5↓) 151.32 0.71 46.8(1.8↓) 334.26 0.67
0.6 81.6(4.6↓) 128.38 0.60 41.8(6.8↓) 309.69 0.61
0.5 79.5(6.7↓) 118.57 0.55 40.1(8.5↓) 288.94 0.57
3.3. Multi-Perspective Self-Correction Engine
Once a step $t_{k}$ is flagged by the AVM, the MSCE is activated to perform a robust and efficient correction. To counteract a model struggles to correct its own mistakes when using a single line of reasoning, the MSCE employs a dual-path correction strategy: Intrinsic Correction (Self-Reflection), the model is presented with the context up to step $s_{(k-1)}$ and its own flawed step $t_{k}$, and is explicitly prompted to review and amend it. Extrinsic Correction (Independent Generation), the model is given the context up to $s_{(k-1)}$ but without the flawed step $t_{k}$. Upon generating two candidate corrections, the MSCE employs a selection mechanism to choose the superior one. We re-apply the Confidence Assessment Score function $Q(\cdot)$ (from Eq. 7, with $U=0$) to both candidates. The candidate with the higher quality score is then integrated into the reasoning chain.
4. Experiment
4.1. Experimental Setup
@@
Models and Datasets We primarily evaluate our method using the LLaMA-3.1-8B-Instruct Ge et al. (2023) and Qwen2.5-Instruct series (3B, 7B, 14B) Team (2024)models. Our evaluation leverages two notably standard math reasoning benchmarks: GSM8K Cobbe et al. (2021), which consists of 8.5K grade school math problems requiring 2-8 reasoning steps, and MATH Hendrycks et al. (2021), a more challenging dataset of 12.5K competition-level math problems. Given the significant computational cost associated with the MATH dataset, we follow recent research practices by evaluating our method on a carefully selected subset, MATH-500 Lightman et al. (2023), which includes a representative sample of 500 problems.
@@
Implementation Details The ASCoT is implemented in PyTorch 1.13.0 and all experiments are conducted on a server with four NVIDIA GeForce RTX 3090 GPUs (24GB). For the IRM module, we use LLMLingua-2 Pan et al. (2024) as the token importance metric to generate compressed CoT data for fine-tuning. To optimize the fine-tuning process, we adopt LoRA Hu et al. (2022) for all fine-tuning procedures, setting the rank $r=8$ and scaling parameter $\alpha=16$ to ensure efficient training. Additional implementation details and configurations can be found in the Appendix.
@@
Evaluation Metrics We evaluate the performance of ASCoT using three primary metrics: Accuracy (%), which represents the percentage of problems for which the final answer is correct. Tokens, the average number of tokens generated in the reasoning chain, measuring computational cost and efficiency during the reasoning process. ActRatio, the actual compression ratio of ASCoT, reflecting the real efficiency of the token compression process.
4.2. Main Results
We evaluate the end-to-end performance of ASCoT against two prevalent CoT length control techniques. The first is Prompt, appends an explicit instruction (e.g., “Please reduce your CoT output by 50%”) to the model’s input, thus guiding the LLM to generate a shorter reasoning trace. The second termed Truncation, imposes a fixed upper bound on the number of output tokens during inference.
The results of LLaMA-3.1-8B-Instruct are summarized in Table 1. As shown, prompt‐based reduction consistently fails to meet the target ratio: even at a nominal 0.5 budget it yields an actual compression ratio of 0.94 on MATH-500, with only marginal savings in CoT tokens. Truncation strictly enforces the specified ratio but at the cost of dramatic performance degradation: at a budget of 0.5, the accuracy on GSM8K plummets by 79.2 points and by 21.2 points on MATH-500, rendering it unsuitable for tasks requiring robust multistep reasoning. In contrast, ASCoT closely matches the specified budget (e.g., ActRatio = 0.79 at Ratio = 0.8 on GSM8K and ActRatio = 0.61 at Ratio = 0.6 on MATH-500) while maintaining robust reasoning performance. In a full budget (Ratio = 1.0), ASCoT marginally outperforms the baseline, achieving 86.9% on GSM8K (+0.7) and 48.8% on MATH-500 (+0.2). In at a 0.5 budget, it maintains respectable accuracies of 79.5% (–6.7) on GSM8K and 40.1% (–8.5) on MATH-500, respectively. These results demonstrate that ASCoT strikes a favorable balance between compression and accuracy, validating its effectiveness.
@@ 4.2. Main Results
Figure 4. Performance of ASCoT on the Qwen2.5-Instruct Series. The results for the 3B, 7B, and 14B models are shown under various compression ratios and compared against the original, uncompressed baseline.

### v4

# Not All Errors Are Created Equal: ASCoT Addresses Late-Stage Fragility in Efficient LLM Reasoning

@@ Abstract
While Chain-of-Thought (CoT) prompting empowers Large Language Models (LLMs), ensuring reasoning reliability remains an open challenge. Contrary to the prevailing cascading failure hypothesis which posits that early errors are most detrimental, we identify a counter-intuitive phenomenon termed Late-Stage Fragility: errors introduced in later reasoning stages are significantly more prone to corrupting final answers. To address this, we introduce ASCoT (Adaptive Self-Correction Chain-of-Thought), a method harmonizing efficiency with robust verification. ASCoT first employs semantic pruning to compress redundant steps, then utilizes an Adaptive Verification Manager (AVM) to prioritize high risk, late-stage steps via a positional impact score, triggering a Multi-Perspective Self-Correction Engine (MSCE) only when necessary. Experiments on GSM8K and MATH-500 demonstrate that ASCoT effectively reallocates computational resources: it reduces token usage by 21%–30% for LLaMA-3.1-8B with negligible accuracy drops ($1.8\%$), achieving a superior trade-off between inference efficiency and reasoning fidelity.
Not All Errors Are Created Equal: ASCoT Addresses Late-Stage Fragility in Efficient LLM Reasoning
[…]
1 Introduction
@@
Figure 1: The top part (a) illustrates the standard CoT reasoning process, which often generates redundant outputs and lacks an effective self-correction mechanism when errors occur. The bottom part (b) depicts the ASCoT method, which incorporates a robust mechanism for self-correction, addressing potential reasoning flaws at each stage.
@@
The advent of CoT has endowed LLMs with remarkable multi-step reasoning capabilities Kojima et al. (2022), enabling them to deconstruct complex problems into a sequence of manageable sub-problems Shinn et al. (2023). Recent studies, such as the o1 from OpenAI  Jaech et al. (2024) and DeepSeekR1  Guo et al. (2025) have shown that extended CoT reasoning improves LLM performance. However, the success of CoT and its variants has introduced challenges: prohibitive computational costs from long reasoning chains Chen et al. (2025) and the inherent brittleness of the reasoning process itself Ke et al. (2025). A single flaw is often sufficient to invalidate an entire process of thought, raising significant concerns about the trustworthiness of these models Pan et al. (2024).
@@
A widely held belief within the research community, often termed the cascading failure hypothesis, posits that the primary threat to CoT integrity lies in early-stage errors Tyen et al. (2023). The logic is intuitive: an initial mistake in problem comprehension or a miscalculation will inevitably propagate through the logical dependency graph of the reasoning chain, contaminating all subsequent steps. In this work, we directly investigate this foundational assumption, seeking to answer a critical question: Are all errors in a CoT created equal in their impact? To find out this, we conduct a systematic error-injection experiment, which introduces controlled symbolic and numeric errors at varying positions within correct reasoning chains. We then instruct the LLM to resume the reasoning process from the point at which the error occurred. Our findings are striking and counter-intuitive. We discover that errors introduced in later stages are more likely to lead to an incorrect final answer. We term this phenomenon Late-Stage Fragility. An early error often appears to trigger the model’s latent self-correction mechanisms, allowing it to recover and still reach the correct solution.
@@
As the model progresses steadily, its reasoning process gains a form of semantic commitment, reducing relevant information entropy. At the later stages, the model is unable to evaluate the validity of the final calculations. This empirical discovery suggests that a significant portion of efforts to secure CoT reasoning may be misdirected. As illustrated in Figure 1, a standard CoT process can be both inefficient due to redundancy and unreliable due to its inability to correct these late-stage errors. Building on a rethinking of how we evaluate and how we correct LLM reasoning, we introduce the ASCoT method. In contrast to standard methods, ASCoT is designed to address both challenges simultaneously, it first employs a semantic pruning stage to enhance efficiency, followed by a targeted, risk-aware correction mechanism that focuses computational resources precisely on the most vulnerable steps to ensure a correct final answer.
Our contributions are summarized as follows:
@@
  • To the best of our knowledge, this is the first to identify and quantify the Late-Stage Fragility phenomenon in CoT reasoning, demonstrating that errors in later steps are significantly more detrimental than those in earlier steps.
@@
  • We introduce ASCoT, a novel method specifically designed to counteract this fragility by precisely identifying and robustly correcting the high-impact, late-stage errors.
@@
  • Our extensive experiments validate the effectiveness of ASCoT. We demonstrate that ASCoT achieves outstanding accuracy on the GSM8k and MATH benchmarks.
2 Related Work
2.1 LLM Reasoning & Efficient Thinking
CoT prompting significantly advances LLM reasoning by decomposing complex problems into intermediate steps Liu et al. (2024b); Team (2024). To mitigate the associated computational costs, recent research explores diverse efficiency-focused strategies, including prompt-based constraints ($e.g.,$ token budgets) Xu et al. (2025); Han et al. (2024), inference-time dynamic early-exit mechanisms Ding et al. (2025), and CoT distillation Wadhwa et al. (2024); Zhang et al. (2026).
2.2 CoT Compression and Pruning
The inherent verbosity of CoT incurs significant computational overhead Wingate et al. (2022) and can induce the overthinking phenomenon, where excessive reasoning introduces errors Kumar et al. (2025). To mitigate this, recent research focuses on compression strategies Xuefei et al. (2023). Approaches vary from token-level pruning based on semantic importance Zhu et al. (2025); Yang et al. (2025) to structural methods like Prune-on-Logic Liu et al. (2025), which filters steps via logic graphs. Notably, even heuristic truncation Wang et al. (2024) has demonstrated efficacy, suggesting that exhaustive reasoning paths are often redundant.
2.3 Self-correction Mechanism
Self-correction mechanisms enhance LLM reliability by iterating through generation, feedback, and refinement phases Kamoi et al. (2024). Current approaches bifurcate based on the feedback source. Intrinsic self-correction relies on the model’s internal critique capabilities without external tools Yang et al. (2025); Zhu et al. (2025). However, these methods often struggle to identify errors due to the scarcity of correction examples in training data Liu et al. (2024a). In contrast, extrinsic self-correction leverages external signals, such as verifier models Kim et al. (2025) or consistency checks, to rigorously assess step-wise correctness.
3 Preliminary
@@
This section establishes the foundational concepts that motivate our work, offering a review of some effective methods for identifying token importance.
@@
Efficient CoT in LLM Reasoning A critical question for CoT efficiency is whether every token in different position contributes equally to the final answer Xia et al. (2025). Several early approaches Hou et al. (2022) measure importance based on the self-information or surprisal from a unidirectional language model:
@@
To address this, more advanced methods like LLMLingua-2 Pan et al. (2024) utilize a bidirectional language model Devlin et al. (2019) ($e.g.,$ BERT-style) to reframe the task as token classification. In this paradigm, the importance of a token $x_{i}$ is defined as the probability that a trained classifier assigns it an essential label ($y_{i}=1$), conditioned on the entire sequence $x$:
@@
where $\theta_{\mathcal{M}{cls}}$ represents the parameters of the trained bidirectional classifier. Such methods confirm that tokens within a CoT sequence are variable in importance. For example, tokens comprising mathematical equations or key entities Ma et al. (2025) are identified as critical, whereas semantic connectors such as “so” and “since” contribute less to the final outcome of the reasoning  Hou et al. (2025). This highlights that CoT efficiency can be improved by pruning these redundant tokens.
Error Propagation in Reasoning Process Beyond efficiency, the primary challenge for CoT is reliability. A single error can invalidate an entire reasoning process Guo et al. (2017). The conventional understanding of this problem is rooted in the cascading failure hypothesis, which posits that errors are most damaging when they occur early in the reasoning chain Jiang et al. (2025). According to this view, an early mistake propagates through the logical dependency graph, contaminating all subsequent steps. In this study, through systematic error-injection experiments, we observe that errors introduced in the later stages of a reasoning chain are significantly more detrimental to the final outcome than identical errors introduced at the beginning.
Figure 2: Illustration of ASCoT Pipeline. ASCoT generates the $CoT_{initial}$ from the target LLM, which is then compressed by IRM module to a predefined ratio $\gamma$. The compressed CoT is verified by AVM module, and if the confidence score exceeds a threshold $\tau$, the problematic steps are sent to MSCE module for error correction. The corrected $CoT_{final}$ is then used for fine-tuning. Finally, ASCoT enables reasoning on new problems with the specified compression ratio $\gamma$.
4 Method
In this section, we detail the architecture of our method. ASCoT generates the $CoT_{initial}$ based on the target LLM. The $CoT_{initial}$ is then processed through the Intelligent Routing Mechanism (IRM) module, reducing the output to a fixed ratio $\gamma$. The remaining CoT is passed through the Adaptive Verification Manager (AVM) module, which performs a rigorous verification process. If the model’s risk score for any CoT step exceeds a predefined threshold $\tau$, indicating a potential error, the corresponding CoT step is forwarded to the Multi-Perspective Self-Correction Engine (MSCE) module. The MSCE module applies internal and external validation techniques to correct errors, resulting in the corrected $CoT_{final}$. Ultimately, the refined, non-redundant $CoT_{final}$ is used for fine-tuning. We illustrate the training and inference process in Figure 2, respectively. We also provide Algorithm 1 for ASCoT.
4.1 Intelligent Routing Mechanism
@@
4.2 Adaptive Verification Manager
@@
Positional Impact Score To quantitatively model the Late-Stage Fragility phenomenon observed in our error-injection experiments (Section 4.3), we introduce the Position Impact Function $I(k)$. This function is an empirical model fitted to our experimental data, designed to capture the escalating impact of errors at later reasoning stages. We normalize this position function as:
@@
where $k$ is the step’s index, $K$ is the total number of steps, $w_{a}$ is a base impact coefficient, and $\alpha$ is a propagation rate. These parameters are determined by fitting the function to our error-injection data. This procedure involves injecting single errors at different steps $k$ in a set of baseline correct reasoning chains and measuring the probability of final answer failure. The function $I(k)$ is then fitted to these empirical data points. Details of the analysis are available in Appendix E.
@@
This score is then compared against a tunable system threshold $\tau$. If $R(t_{k})>\tau$, the step is flagged as high risk and passed to the MSCE.
4.3 Multi-Perspective Self-Correction Engine
Once a step $t_{k}$ is flagged by the AVM, the MSCE is activated to perform a robust and efficient correction. To counteract a model struggles to correct its own mistakes when using a single line of reasoning, the MSCE employs a dual-path correction strategy: Intrinsic Correction, the model is presented with the context up to step $s_{(k-1)}$ and its own flawed step $t_{k}$, and is explicitly prompted to review and amend it. Extrinsic Correction, the model is given the context up to $s_{(k-1)}$ but without the flawed step $t_{k}$. Upon generating two candidate corrections, the MSCE employs a selection mechanism to choose the superior one. We re-apply the Confidence Assessment Score function $Q(\cdot)$ (from Eq. 7, with $U=0$) to both candidates. The candidate with the higher quality score is then integrated into the reasoning chain.
Figure 3: The details of verification and self-correction. AVM and MSCE working together to correct a faulty step $t_{k}$. AVM computes $R(t_{k})$. MSCE then applies intrinsic and extrinsic correction to produce the final corrected $CoT_{final}$.
Table 1: Experimental results of ASCoT on LLaMA-3.1-8B-Instruct. We report Accuracy, average CoT token count (Tokens) and actual compression ratio (ActRatio) for comparison. The compression ratio $\gamma$ is set to 0.51.00.5-1.0.
@@ 4.3 Multi-Perspective Self-Correction Engine
Original 1.0 86.2($0\downarrow$) 213.17 1.00 48.6($0\downarrow$) 502.60 1.00
[…]
5 Experiment
5.1 Experimental Setup
@@
Models and Datasets We primarily evaluate our method using the LLaMA-3.1-8B-Instruct Dubey et al. (2024) and Qwen2.5-Instruct series (3B, 7B, 14B) Team (2024)models. Our evaluation leverages two notably standard math reasoning benchmarks: GSM8K Cobbe et al. (2021), which consists of 8.5K grade school math problems requiring 2-8 reasoning steps, and MATH Hendrycks et al. (2021), a more challenging dataset of 12.5K competition-level math problems. Given the significant computational cost associated with the MATH dataset, we follow recent research practices by evaluating our method on a carefully selected subset, MATH-500 Lightman et al. (2023), which includes a representative sample of 500 problems.
@@
Implementation Details The ASCoT is implemented in PyTorch 1.13.0 and all experiments are conducted on a server with four NVIDIA GeForce RTX 3090 GPUs (24GB). For the IRM module, we use LLMLingua-2 Pan et al. (2024) as the token importance metric to generate compressed CoT data for fine-tuning. To optimize the fine-tuning process, we adopt LoRA Hu et al. (2022) for all fine-tuning procedures, setting the rank $r=8$ and scaling parameter $\alpha=16$ to ensure efficient training. Additional implementation details and configurations can be found in the Appendix D.
@@
Evaluation Metrics We evaluate the performance of ASCoT using three primary metrics: Accuracy (%), which represents the percentage of problems for which the final answer is correct. Tokens, the average number of tokens generated in the reasoning chain, measuring computational cost and efficiency during the reasoning process. ActRatio, the actual compression ratio of ASCoT, reflecting the real efficiency of the token compression.
@@
Figure 4: Performance of ASCoT on the Qwen2.5-Instruct Series. The results for the 3B, 7B, and 14B models are shown under various compression ratios and compared against the original, uncompressed baseline.
5.2 Main Results
We benchmark ASCoT against two standard length-control baselines: instruction-based Prompting and hard Truncation. The comparative results on LLaMA-3.1-8B are presented in Table 1.
Comparison with Baselines Existing methods struggle to balance compression controllability with reasoning fidelity. Prompting fails to strictly adhere to token budgets. Truncation also incurs catastrophic performance degradation when it enforces strict length constraints.
In contrast, ASCoT demonstrates precise controllability and robust performance. It aligns closely with target ratios (e.g., ActRatio 0.79 at a 0.8 target) while preserving reasoning accuracy. Notably, at a full budget ($\gamma=1.0$), ASCoT marginally outperforms the original baseline (+0.7% on GSM8K), suggesting that our verification mechanism effectively filters noise. Even under aggressive compression ($\gamma=0.5$), ASCoT maintains a respectable 79.5% accuracy on GSM8K, significantly outperforming truncation based methods.
Scalability Analysis Table 2 details the performance across the Qwen2.5-Instruct series (3B, 7B, 14B) to evaluate scalability. ASCoT achieves substantial token reduction with minimal accuracy loss across all scales. For the 3B model, a 0.9 compression ratio reduces token usage by approximately 16% with only a 0.5 point accuracy drop. This resilience becomes more pronounced as model size increases. The 14B model exhibits exceptional robustness. Even when halving the token budget, it retains 91.6% accuracy (a mere 1.5% drop). This trend suggests a positive correlation between model scale and reasoning redundancy, larger models tend to generate more steps that ASCoT effectively identifies and prunes without compromising the final answer, demonstrating its potential to significantly lower inference costs for large-scale deployments without sacrificing reliability.